Using Taylor series expression, find a power series solution for the equation

2x2d2y/dx2-xdy/dx+(x-5)y=0 in powers of (x-1) if y(1)=4 and y'(1)=2


1
Expert's answer
2021-08-24T08:32:25-0400

Solution;

Taylors series expansion of f(x) at x=a is given as follows;

"f(x)=f(a)+f'(a)(x-a)+\frac{f''(a)}{2!}(x-a)^2+\frac{f'''(a)}{3!}(x-a)^3+..."

Given;x=1

We have;

"f(x)=f(1)+f'(1)(x-1)+\frac{f''(1)}{2!}(x-1)^2+\frac{f'''(1)}{3!}(x-1)+..."

Since;

y(1)=4

y'(1)=2

Also;

"2x^2\frac{d^2y}{dx^2}-x\frac{dy}{dx}+(x-5)y=0"

By direct substitution;

"2(1^2)y''-1(2)+(1-5)(4)=0"

"2y''-2-16=0"

"y''(1)=9"

Now we find the triple derivative at x=1;

"2x^2\frac{d^2y}{dx^2}-x\frac{dy}{dx}+(x-5)y=0"

Factorise to obtain;

"2x^2\frac{d^2y}{dx^2}-x\frac{dy}{dx}+xy-5y=0"

Differentiate by applying the product rule;

"[2x^2\frac{d^3y}{dx^3}+\frac{d^2y}{dx^2}(4x)]-[x\frac{d^2y}{dx^2}+\frac{dy}{dx}]+[x\frac{dy}{dx}+y]-0=0"

By direct substitution;

"2(1^2)y'''+(9)(4\u00d71)-(1)(9)-2+(1)(2)+4=0"

Simplify to obtaining;

"y'''(1)=-\frac{31}{2}"

Hence we have the power solution series as;

"f(x)=4+2(x-1)+\frac9{2!}(x-1)^2-\frac{31}{2×3!}(x-1)^3+..."

Simplify;

"f(x)=4+2(x-1)+\frac92(x-1)^2-\frac{31}{12}(x-1)^3+..."